3.4.14 Perform the indicated calculation. 5P2 10P4 5P2 10P4 (Round to four decimal places as needed.)
Answer #1

Answer: n! Required formula : In Pop (n-r! 5! 51 5P2 = = (5-2)! 3! 11 5x4x8! ZI = 20 겨 10? 10 P4 a 10! (10-4)! II 6!- 10x9x8x1x 2 8 = 5040 Now, 5 P_ - 20 5040 10 P4 - 0.00396825 39+ 0.0040 ।

Your rating is important

Please rate (like), thank you.